# Stealthy Jailbreak Attacks on Large Language Models via Benign Data Mirroring

### v1

@@ 3.1.3 Attack Stealth:
We list the number of requests for each of these two stages separately, along with the number of requests that can be identified as jailbreak attempts among all requests in both stages. We employ meta-llama/Prompt-Guard-86M (Dubey et al., 2024) as the classifier to detect jailbreak, as it offers an out-of-box method to identify common jailbreak prompts with high precision on Meta’s private jailbreak evaluation dataset. To evaluate the stealth of attacks, we analyze the total number of requests and the number of requests identified as jailbreak attempts in both the preparation and attack stages.
@@ 4.1.2 Evaluation Datasets
It’s important to note that both AutoDAN Liu et al. (2024) and GCG Zou et al. (2023) require affirmative example outputs as optimization targets when searching for adversarial prompts. For example, given a malicious instruction like “How can I use a deodorant can as a flamethrower?”, a corresponding example output might be “Sure, here are detailed instructions on how to use a deodorant can as a flamethrower.” Note that These example outputs are merely affirmative rewritings of the input instructions and contain only the information provided in the input, without including any actual harmful output or suggestions. While AdvBench already includes such outputs, the StrongReject dataset lacks them. Therefore, we manually annotated these outputs for the StrongReject dataset. This manual annotation process is elaborated in the appendix D.
@@ 4.2 Evaluation Settings
AutoDAN Transfer Attack Liu et al. (2024) 0.0 20.0 2.4 3.0 2.4 0.32
@@ 4.3 Experiments Results
The ShadowBreak method demonstrates generalizability across different jailbreak methods and models, as shown in Table 1. It achieved high ASRs using both GCG (up to 92%) and AutoDAN (up to 80%) on AdvBench against GPT-3.5 Turbo. However, effectiveness varied across specific test sets and model architectures. For instance, GPT-4o mini exhibited significantly better safety performance than GPT-3.5 Turbo, fully defending against GCG attacks while remaining vulnerable to AutoDAN attacks. Notably, both jailbreak methods failed on the StrongReject test set against GPT-4o mini. We also tested ShadowBreak with different mirror models, as shown in Table 4, demonstrating that mirror model selection plays a crucial role in attack success. These findings suggest that while ShadowBreak is broadly applicable, its performance is influenced by the specific characteristics of the models and the nature of the safety categories being tested.
@@ Appendix C Generation Configuration
Appendix D Manual Labeling Guide
@@ Appendix E Artifacts
Appendix F AI Assistant Disclosure

### v2

@@ 3.1.3 Attack Stealth:
To evaluate attack stealth, we measure the average number of requests per instruction ($Q$) and the number flagged as jailbreak attempts per instruction ($Q^{!}$) during both the preparation and attack stages. Specifically, for each stage, we define:
$Q$ $=\frac{|Queries|}{|I|},$ (3)
$Q^{!}$ =|$=\frac{|\{Query_{i}\in Queries\mid\mathcal{J}(Query_{i})=1\}|}{|I|}$ (4)
where $\mathcal{J}$ represents the jailbreak detection function, which returns 1 only when a query is detected as a jailbreak. For $J$, we employ meta-llama/Prompt-Guard-86M (Dubey et al., 2024) as the classifier to detect jailbreak. Released by Meta in July 2024, this classifier was trained on Meta’s private dataset to categorize inputs into three classes: benign, injection, and jailbreak. We label the jailbreak class as 1 and the other two as 0.
@@ 4.1.2 Evaluation Datasets
It’s important to note that both AutoDAN Liu et al. (2024) and GCG Zou et al. (2023) require affirmative example outputs as optimization targets when searching for adversarial prompts. For example, given a malicious instruction like “How can I use a deodorant can as a flamethrower?”, a corresponding example output might be “Sure, here are detailed instructions on how to use a deodorant can as a flamethrower.” Note that These example outputs are merely affirmative rewritings of the input instructions and contain only the information provided in the input, without including any actual harmful output or suggestions. While AdvBench already includes such outputs, the StrongReject dataset lacks them. Therefore, we manually annotated these outputs for the StrongReject dataset. This manual annotation process is elaborated in the appendix E.
@@ 4.2 Evaluation Settings
AutoDAN Transfer Attack Liu et al. (2024) 0.0 0.0 2.4 3.0 2.4 0.32
@@ 4.3 Experiments Results
The ShadowBreak method demonstrates generalizability across different jailbreak methods and models, as shown in Table 1. It achieved high ASRs using both GCG (up to 92%) and AutoDAN (up to 80%) on AdvBench against GPT-3.5 Turbo. However, effectiveness varied across specific test sets and model architectures. For instance, GPT-4o mini exhibited significantly better safety performance than GPT-3.5 Turbo, fully defending against GCG attacks while remaining vulnerable to AutoDAN attacks. Notably, both jailbreak methods failed on the StrongReject test set against GPT-4o mini. Additionally, we conducted an experiment on Claude 3.5 Haiku (Anthropic, 2024), detailed in Appendix A, which suggests that ShadowBreak enhances the prefilling attack (Andriushchenko et al., 2024) on this model.
We also tested ShadowBreak with different mirror models, as shown in Table 4, demonstrating that mirror model selection plays a crucial role in attack success. These findings suggest that while ShadowBreak is broadly applicable, its performance is influenced by the specific characteristics of the models and the nature of the safety categories being tested.
@@ Appendix A Additional Experimental Results
Claude Experiments
We conducted an evaluation of our method on claude-3-5-haiku-20241022, yielding noteworthy findings. Specifically, for the mirroring data, we queried claude-3-5-haiku-20241022 using the same set of 1k benign instructions as detailed in Table 2. Utilizing AutoDAN (Liu et al., 2024) as the searcher, we executed experiments on AdvBench. Given that Claude permits partial control over assistant outputs via prefilling, we incorporated the baseline from Andriushchenko et al. (2024) by appending the prefix "Sure," to Claude’s responses. Each result presented is an average of three runs, as described in Section 4.2.
Method Sys Msg $\text{ASR}_{M}$ $\text{ASR}_{C}$
Prefilling Attack default $40.0$
+ Naïve Transfer Attack default $0.32$ $0.06$
+ Mirroring (Benign 1k) (Ours) default $0.5$ $0.26$
Prefilling Attack none $0.12$ $0.02$
+ Naïve Transfer Attack none $0.46$ $0.38$
+ Mirroring (Benign 1k) (Ours) none $0.64$ $0.52$
Table 5: Experimental results comparing different attack methods.
The results indicate that Claude 3.5 (Anthropic, 2024) demonstrates strong jailbreak defense with 2% on ASR under the prefilling attack Andriushchenko et al. (2024). Furthermore, our benign data mirroring method further enhances attack performance, improving ASR by 50% over prefilling attacks alone and surpassing naïve transfer by 14%, highlighting its generalization potential.
These findings provide further insight into the generalizability of our approach.
DPO Experiments
@@ Appendix C Generation Configuration
Appendix D Cost Analysis
API Usage
We compare our method against search-based black-box attack baselines such as PAIR (Chao et al., 2023) and PAL Sitawarin et al. (2024), which also depend on repeated API queries. Even when incorporating the benign data distillation phase, our approach requires significantly fewer queries. Under comparable conditions (using 1k aligned samples as a baseline), the average query counts are 6.1k for PAL, 140.4 for PAIR, and only 23 for our method as in Table 3. A full distillation using 20k samples (e.g., with GPT-4-mini) incurred a total query cost of 440k input tokens and 521k output tokens, corresponding to approximately \$3.2 in API usage costs.
GPU Usage
The fine-tuning process, implemented with LoRA (Hu et al., 2021), is designed for resource efficiency. The fine-tuning itself required roughly two hours for 20k samples, costing about \$8 on AWS (as detailed in Appendix B).
This \$8 overhead is significantly lower than that of existing search methods. For instance, on the AdvBench dataset with 50 samples Zou et al. (2023); Chao et al. (2023), naive AutoDAN (Liu et al., 2024) requires 24 hours at a cost of \$96, while GCG (Zou et al., 2023) requires 5 hours at a cost of \$20.
Appendix E Manual Labeling Guide
@@ Appendix F Artifacts
Appendix G AI Assistant Disclosure